OpenStudy (anonymous):

how do i find the antiderivative of 1 / √x

OpenStudy (anonymous):

\[\int \frac{1}{\sqrt{x}}dx => \int \frac{1}{(x)^{1/2}}dx => \int x^{-1/2}dx\]

OpenStudy (anonymous):

well apparently the answer is 2√x - 10. which makes no sense.

OpenStudy (anonymous):

well have you tried it

OpenStudy (anonymous):

and it cannot be \(2\sqrt{x}-10\)